# Stochastic Gradient Descent Jittering for Inverse Problems: Alleviating the Accuracy-Robustness Tradeoff

### v1

@@ Introduction
$={\bm{x}}_{2}^{\prime}+\eta\big{(}f_{\theta}({\bm{x}}_{1}^{\prime})-f_{\theta}({\bm{x}}_{1}^{sgd})\big{)}+\eta(1-\eta)\big{(}f_{\theta}({\bm{x}}_{0}^{\prime})-f_{\theta}({\bm{x}}_{0}^{sgd})\big{)}$
$\qquad-\big{(}\eta(1-\eta)^{2}{\bm{w}}_{0}+\eta(1-\eta){\bm{w}}_{1}+\eta{\bm{w}}_{2}\big{)}$
@@
$={\bm{x}}_{k+1}^{\prime}+\sum_{i=0}^{k}\eta(1-\eta)^{k-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))$
$\qquad-\sum_{i=0}^{k+1}\eta(1-\eta)^{k+1-i}{\bm{w}}_{i}.$
@@
$\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{sgd}||_{2}^{2}=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}~{}||\sum_{i=0}^{K}\eta(1-\eta)^{K-i}{\bm{w}}_{i}||_{2}^{2}$
$+\mathbb{E}~{}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))||_{2}^{2}$
$-2\mathbb{E}\left[\langle\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}f_{\theta}({\bm{x}}_{i}^{sgd}),\sum_{i=0}^{K}\eta(1-\eta)^{K-i}{\bm{w}}_{i}\rangle\right].$
@@
In addition, we can expand the regularization term in SGD training via Taylor expansion. Let $\bm{\delta}_{k+1}=\sum_{i=0}^{k}\eta(1-\eta)^{k-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))+\sum_{i=0}^{k+1}\eta(1-\eta)^{k+1-i}{\bm{w}}_{i}$. Assumption 3 ensures the existence of the gradient and the Hessian of $f_{\theta}$. We have,
$f_{\theta}({\bm{x}}_{k+1}^{sgd})\approx f_{\theta}({\bm{x}}_{k+1}^{\prime})+\bm{\delta}_{k+1}^{\top}\nabla_{\bm{x}}f_{\theta}({\bm{x}}_{k+1}^{\prime})+\frac{1}{2}\bm{\delta}_{k+1}^{\top}\nabla^{2}_{\bm{x}}f_{\theta}({\bm{x}}_{k+1}^{\prime})\bm{\delta}_{k+1}.$
The higher the iteration $k$, the larger penalty weight $\eta(1-\eta)^{K-1-i}$ becomes. Thus with a large $K$, SGD jittering promotes a smoother and wider landscape for the iterative inputs of the $f_{\theta}$. We also refer to this as a flat minimum because the update of ${\bm{x}}_{k}$ solves the lower-level minimization problem. This is consistent with the observations in (Lim et al. 2021), which demonstrate that adding random Gaussian noise to recurrent neural networks layer-wise promotes a small Hessian in the last single-block neural network in terms of its inputs. While it is widely observed that flat minima in the loss landscape (with respect to network parameters) imply better generalization as summarized in Related Works, (Lim et al. 2021) does not clarify why flat minima with respect to hidden-layer inputs improve generalization in regression settings. Our work fills this gap in the context of solving IPs. Now, we state the robustness result,
Theorem 3.
$R_{e}(\theta)=\mathbb{E}_{({\bm{x}},{\bm{y}})\in\mathcal{D},{\bm{e}}\in\mathcal{P}_{e}}~{}\left[||{\bm{x}}-H_{\theta}({\bm{y}}+{\bm{e}})||^{2}_{2}\right].$ (9)
$R_{e}(\theta^{sgd})\leq R_{e}(\theta^{mse}).$
The proof is similar to the one in proving generalization, we include the full proof in the Appendix. The gradient and Hessian of $f_{\theta}$ together help characterize the generalization and robustness of MBAs.
Figure 2: Seismic deconvolution under adversarial attack.
Experiments and Discussion
In our experiments, we use a toy denoising example to validate our theorems and demonstrate the proposed training scheme using seismic deconvolution and magnetic resonance imaging (MRI) reconstruction. For each task, we train the same LU architecture using different training schemes. We then evaluate the models using in-distribution testing data, adversarial attacks with various strengths, and task-dependent OOD data. Detailed training procedures are provided in the Appendix.
Toy problem: 2-dimensional Point Denoising
To validate the Theorems, we compare the robustness and generalization using regular MSE, AT and SGD-jittering in a point denoising problem. Assume ${\bm{x}}=(0,0)^{\top}$, we generate noisy ${\bm{y}}$ by adding random zero-mean Gaussian noise with a variance of 0.01. We use 200 samples for training and 50 for evaluation. Figure 1 shows the noisy points ${\bm{y}}$ and the denoised results $\hat{{\bm{x}}}$ (blue) across four evaluation modes, with the ground-truth in red. In-distribution data are sampled from the same distribution as the training data, centered at $(0,0)^{\top}$. The worst-case attack finds a vector within a 0.01 $\ell_{2}$ distance around ${\bm{y}}$ that maximizes the squared error, while the average-case attack uniformly samples attack vectors within the same ball and adds to ${\bm{y}}$. SGD jittering and AT are more robust than MSE training in both metrics. AT aims to promote worst-case robustness, which implies average-case robustness for attacks within the same $\ell_{2}$ ball around ${\bm{y}}$, though its average-case performance may not be optimal. We observe that SGD jittering has even tighter clusters in its reconstructions under average-case attacks than AT. Finally, we evaluate generalization accuracy by reconstructing the slightly biased ground truth $(0.005,0)^{\top}$ from its corresponding measurements. With the introduction of bias, SGD jittering achieves a more accurate reconstruction compared to the new ground truth. Notice that MSE training underperforms on in-distribution data compared to SGD jittering, likely because the noise injection in SGD jittering helps the model escape from local minima in training.
Figure 3: Seismic deconvolution with an additional layer.
Figure 4: MRI evaluation (a) under adversarial attack obtained from projected gradient descent with norm no more than $\epsilon=1$. Each row represents an MRI reconstruction of a sample from fastMRI (Zbontar et al. 2018a). (b) MRI evaluation using OOD tumor knee data from (Bickle and Jin 2021).
Seismic Deconvolution
Seismic deconvolution is a crucial problem in geophysics. It aims to reverse the convolution effects in the received signal, which occurs when artificial source waves travel through the Earth’s layers. The goal is to extract sparse reflectivity series from recorded seismic data. The training data is generated following the same procedure as in (Iqbal et al. 2019; Guan et al. 2024). To demonstrate the generalization accuracy, we generate OOD testing data by adding an arbitrary horizontal layer with magnitudes of 0.1 and 0.3 to the ground truth, reflected accordingly in the measurements. Table 1 numerically compares the average testing peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) of the reconstructions. Notice that for OOD data, the numerical values are very similar, as the additional layer has a small magnitude relative to the overall measurement. However, the difference can be clearly observed in the figures. Figure 2 shows two deconvolution examples for each trained model under adversarial attacks. MSE training produces artifacts, while both AT and SGD jittering yield results closer to the ground truth. Notably, in the top example, AT fails to recover the small horizontal layer in the center of the boxed area, but SGD jittering successfully captures it. Figure 3 demonstrates the generalization ability with two different magnitude levels. For the OOD example with the larger magnitude, both SGD jittering and MSE training recover a cleaner layer compared to AT. More notably, with the smaller magnitude, AT perceives the line as an unwanted perturbation and barely recovers it, whereas the result from SGD jittering is more distinctly visible.
PSNR / SSIM Test data Adv. Attack OOD small OOD large
MSE training 34.64 / 0.921 28.85 / 0.829 34.57 / 0.918 34.57 / 0.917
AT 33.50 / 0.903 30.27 / 0.849 33.45 / 0.902 33.43 / 0.902
Input Jittering 32.92 / 0.882 30.10 / 0.832 32.91 / 0.882 32.90 / 0.882
SGD jittering 35.10 / 0.928 29.89 / 0.842 34.93 / 0.927 34.93 / 0.927
Table 1: Seismic deconvolution evaluation for in-distribution data (column 2), under adversarial attack with $\epsilon=1$ (column 3), and for OOD data with small and large magnitudes (columns 4 and 5). The best and second-best performances are in bold and underlined respectively.
Accelerated MRI Reconstruction
Accelerated MRI aims to recover human-interpreted body structures from partial k-space measurements. We train models with single-coil knee MRI from the fastMRI dataset (Zbontar et al. 2018b) with 4×4$\times$4 × acceleration, or $1/4$ of the measurements in k-space is used for reconstruction. To assess generalization, we use a different knee dataset from Bickle and Jin (2021), which includes giant tumor cells not present in the training dataset. Table 2 numerically compares the average testing PSNR and SSIM of the reconstructions for in-distribution fastMRI data, under adversarial attack with $\epsilon=1$ and OOD tumor cell knee data. Our findings indicate that MSE training and SGD jittering exhibit comparable performance on in-distribution data, but SGD jittering demonstrates greater robustness and superior performance on OOD data. Adversarial training (AT) yields the most robust model under attack, though it comes with a significant tradeoff in performance for both in-distribution and OOD data. Additionally, as shown in Figure 4, reconstructions using AT appear blurrier than others, a similar observation noted in black-box IP solvers by Krainovic, Soltanolkotabi, and Heckel (2023).
PSNR / SSIM fastMRI Adv. Attack Tumor Cell
MSE training 28.21 / 0.603 25.68 / 0.382 29.92 / 0.779
AT 27.68 / 0.564 27.17 / 0.549 27.74 / 0.597
Input Jittering 28.18 / 0.595 25.05 / 0.420 29.97 / 0.740
SGD jittering 28.22 / 0.607 26.77 / 0.552 30.36 / 0.788
Conclusion
Robustness and generalization accuracy are both crucial in solving IPs, yet current analyses often focus on simple black-box solvers that enhance robustness at the cost of accuracy. While empirical studies have examined the robustness of MBAs, there is a lack of theoretical analysis. In this work, we introduce a novel and easily implemented SGD jittering training scheme to address the tradeoff between robustness and accuracy in solving IPs. Our method proves effective for high-quality model-based IP solvers. Through mathematical analysis, we demonstrate that this approach implicitly regularizes the gradient and Hessian of the neural network with respect to the input, resulting in improved generalization and enhanced robustness compared to traditional MSE training. Experimental results show that while MSE training is vulnerable to adversarial attacks and AT recovers smooth estimates, SGD jittering consistently produces robust, high-quality outcomes.
Appendix A Appendix
Appendix B Notations
We clarify the notations used in the paper.
Appendix C Proof of Lemma 1
Lemma 4.
(Restate Lemma 1. Convergence of MBAs-SGD for IPs) Let $F({\bm{x}})=\frac{1}{2}||{\bm{y}}-\bm{A}{\bm{x}}||_{2}^{2}+r({\bm{x}})$ denote the lower-level objective function. Assume $f_{\theta}({\bm{x}})$ is $L$-Lipschitz continuous,
$||f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}\leq L||\Delta||_{2},\quad\forall{\bm{x}},\Delta$
With assumption 1, and step-sizes $\eta_{k}<2/L$, SGD in minimizing $F({\bm{x}})$ converges to a stationary point.
Proof.
The lemma can be easily derived from the Theorem in (Ghadimi and Lan 2013), which assumes $\nabla_{\bm{x}}F({\bm{x}})$ is $L$-Lipschitz continuous.
$||\nabla_{\bm{x}}F({\bm{x}}+\Delta)-\nabla_{\bm{x}}F({\bm{x}})||_{2}$
$=$ $||\bm{A}^{\top}(\bm{A}({\bm{x}}+\Delta)-{\bm{y}})+f_{\theta}({\bm{x}}+\Delta)-\bm{A}^{\top}(\bm{A}{\bm{x}}-{\bm{y}})-f_{\theta}({\bm{x}})||_{2}$
$=$ $||\bm{A}^{\top}\bm{A}\Delta+f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}$
$\leq$ $||\bm{A}^{\top}\bm{A}||\,||\Delta||_{2}+||f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}$
$\leq$ $\max_{i=1,..,n}\lambda_{i}||\Delta||_{2}+||f_{\theta}({\bm{x}}+\Delta)-f_{\theta}({\bm{x}})||_{2}$
where $\lambda_{i}$ be the eigenvalues of $\bm{A}^{\top}\bm{A}$, and let $\mu=\max_{i=1,..,n}\lambda_{i}$. Since $\nabla F$ is $L$-Lipshitz, we show that for MBAs-SGD in solving denoising problems, it is equivalent to $(L-\mu)$-Lipschitz continuous of $f_{\theta}$. With the same assumptions on step sizes and the SGD jittering noise, we finish the proof. ∎
Appendix D Proof of Theorem 3 - robustness
Theorem 5.
@@ Appendix D Proof of Theorem 3 - robustness
(Restate Theorem 3) Let ${\bm{e}}$ be the perturbation vector and let $\mathcal{P}_{e}$ be iid zero-mean Gaussian distribution such that $\mathbb{E}||{\bm{e}}||_{2}^{2}=\sigma_{\bm{e}}^{2}$. The average-case robustness risk for denoising problem is as follows,
$R_{e}(\theta)=\mathbb{E}_{({\bm{x}},{\bm{y}})\in\mathcal{D},{\bm{e}}}~{}\left[||{\bm{x}}-H_{\theta}({\bm{y}}+{\bm{e}})||^{2}_{2}\right].$ (10)
@@
$=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))||^{2}_{2}$
$\quad+\mathbb{E}||{\bm{e}}||_{2}^{2}+2\mathbb{E}\left[<{\bm{e}},\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))>\right]$
$=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))||^{2}_{2}$
$\quad+\mathbb{E}||{\bm{e}}||_{2}^{2}+2\mathbb{E}\left[<{\bm{e}},-\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}-f_{\theta}({\bm{x}}_{i}^{at})>\right]$
$=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))||^{2}_{2}+\sigma_{\bm{e}}^{2}$
@@
$\quad+2\mathbb{E}\left[\langle{\bm{e}},-\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}-f_{\theta}({\bm{x}}_{i}^{at})\rangle\right]$
@@
The first term is the mean-squared testing loss for in-distribution data. The second term is an additional penalty that appears in SGD jittering training. When picking $\sigma_{{\bm{w}}_{k}}^{2}$ for all $k$ and $\sigma_{\bm{e}}^{2}$ such that $\sum_{k=0}^{k}\eta^{2}(1-\eta)^{2(k-i)}\sigma_{w_{i}}^{2}=\sigma_{\bm{e}}^{2}$, this term in robustness risk is a special case in SGD jittering risk. The last term also appears in SGD jittering risk, when the above condition for jittering level and robustness level are satisfied, ${\bm{e}}$ and $\sum_{i=0}^{K}\eta(1-\eta)^{K-i}{\bm{w}}_{i}$ are from the same distribution. Therefore, for models that perform equally well in mean-squared error for in-distribution testing data, minimizing SGD jittering risk minimizes overall robustness. ∎
Appendix E Training Details

### v2

@@ Appendix A Notations
$={\bm{x}}_{2}^{\prime}+\eta\big{(}f_{\theta}({\bm{x}}_{1}^{\prime})-f_{\theta}({\bm{x}}_{1}^{sgd})\big{)}+\eta(1-\eta)\big{(}f_{\theta}({\bm{x}}_{0}^{\prime})-f_{\theta}({\bm{x}}_{0}^{sgd})\big{)}-\big{(}\eta(1-\eta)^{2}{\bm{w}}_{0}+\eta(1-\eta){\bm{w}}_{1}+\eta{\bm{w}}_{2}\big{)}$
@@
$={\bm{x}}_{k+1}^{\prime}+\sum_{i=0}^{k}\eta(1-\eta)^{k-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))-\sum_{i=0}^{k+1}\eta(1-\eta)^{k+1-i}{\bm{w}}_{i}.$
@@
$\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{sgd}||_{2}^{2}$ $=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}~{}||\sum_{i=0}^{K}\eta(1-\eta)^{K-i}{\bm{w}}_{i}||_{2}^{2}+\mathbb{E}~{}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd}))||_{2}^{2}$
$\quad+2\mathbb{E}~{}\left[\langle\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{sgd})),\sum_{i=0}^{K}\eta(1-\eta)^{K-i}{\bm{w}}_{i}\rangle\right].$
@@
Appendix D Proof of Robustness (Theorem 7.5)
Theorem D.1.
@@
(Restate Theorem 7.5) Let ${\bm{e}}$ be the perturbation vector and let $\mathcal{P}_{e}$ be iid zero-mean Gaussian distribution such that $\mathbb{E}||{\bm{e}}||_{2}^{2}=\sigma_{\bm{e}}^{2}$. The average-case robustness risk for denoising problem is as follows,
$R_{e}(\theta)=\mathbb{E}_{({\bm{x}},{\bm{y}})\sim\mathcal{D},{\bm{e}}\sim\mathcal{P}_{e}}~{}\left[||{\bm{x}}-H_{\theta}({\bm{y}}+{\bm{e}})||^{2}_{2}\right].$ (14)
@@
$=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))||^{2}_{2}+\mathbb{E}||{\bm{e}}||_{2}^{2}+2\mathbb{E}\left[\langle{\bm{e}},\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))\rangle\right]$
@@
$=\mathbb{E}||{\bm{x}}-{\bm{x}}_{K}^{\prime}||_{2}^{2}+\mathbb{E}||\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))||^{2}_{2}+\sigma_{\bm{e}}^{2}+2\mathbb{E}\left[\langle{\bm{e}},\sum_{i=0}^{K-1}\eta(1-\eta)^{K-1-i}(f_{\theta}({\bm{x}}_{i}^{\prime})-f_{\theta}({\bm{x}}_{i}^{at}))\rangle\right]$
@@
The first term is the mean-squared testing loss for in-distribution data. The second term is an additional penalty that appears in SGD jittering training. When picking $\sigma_{{\bm{w}}_{k}}^{2}$ for all $k$ and $\sigma_{\bm{e}}^{2}$ such that $\sum_{k=0}^{k}\eta^{2}(1-\eta)^{2(k-i)}\sigma_{w_{i}}^{2}=\sigma_{\bm{e}}^{2}$, this term in robustness risk is a special case in SGD jittering risk. The last term also appears in SGD jittering risk, when the above condition for jittering level and robustness level are satisfied, ${\bm{e}}$ and $\sum_{i=0}^{K}\eta(1-\eta)^{K-i}{\bm{w}}_{i}$ are from the same distribution. Therefore, for models that perform equally well in MSE for in-distribution testing data, minimizing SGD jittering risk minimizes overall robustness. ∎
Appendix E Extra Stochastic Proximal Gradient Descent Experiments
Similar to gradient descent variants of loop unrolling architectures, for proximal gradient descent (PGD) variants, we train a PGD LU with regular MSE loss. Then finding the worst-case attack to ${\bm{y}}$ over all PGD updates.
E.1 Seismic Deconvolution
In the same seismic deconvolution problem setup discussed in the main manuscript, but now solved using proximal gradient variants, Figure 7(a) shows the performance of the trained models under an adversarial attack with $\epsilon=1$. Figure 7(b) focuses on reconstructing the additional layer highlighted in red boxes. Table 3 provided numerical evaluations. The proposed SPGD jittering demonstrates greater robustness compared to MSE training while better preserving detailed information than adversarial training (AT).
Figure 7: Visualization of seismic deconvolution using (stochastic) proximal gradient descent variants within loop unrolling architectures. Figure (a) illustrates the reconstruction performance under adversarial attack with $\epsilon=1$. Figure (b) evaluates the model’s performance on out-of-distribution (OOD) seismic data, where the ground-truth reflectivity includes an additional small horizontal layer, highlighted in the red boxes.
PSNR / SSIM in-distribution Adv. Attack OOD
MSE training 34.91 / 0.923 28.78 / 0.842 34.810 / 0.920
AT 32.99 / 0.899 31.44 / 0.875 32.966 / 0.899
Input Jittering 33.02 / 0.899 30.56 / 0.861 32.910 / 0.895
SPGD Jittering 35.05 / 0.931 29.93 / 0.870 34.952 / 0.929
Table 3: PGD variants of LU are evaluated on seismic deconvolution across three scenarios: in-distribution data (column 2), adversarial attacks (column 3), and out-of-distribution (OOD) data (column 4). The best performance in each case is highlighted in bold, while the second-best performance is underlined.
E.2 Accelerated MRI Reconstruction
We complete the missing numerical results discussed in the main manuscript for 4×4$\times$4 × accelerated MRI reconstruction using SPGD jittering in Table 4.
PSNR / SSIM fastMRI Adv. Attack Tumor Cell
MSE training 27.93 / 0.582 22.82 / 0.179 29.39 / 0.771
AT 27.20 / 0.533 26.78 / 0.518 27.13 / 0.560
Input Jittering 24.88 / 0.299 24.44 / 0.297 25.85 / 0.478
SPGD Jittering 28.15 / 0.597 27.56 / 0.577 29.762 / 0.777
Table 4: PGD variants of LU are evaluated on MRI data across three scenarios: in-distribution data (column 2), adversarial attacks (column 3), and out-of-distribution (OOD) data (column 4). The best performance in each case is highlighted in bold, while the second-best performance is underlined.
Appendix F Compared to Other Baselines for MRI reconstruction
We evaluate our method against a widely used robustness technique Lipschitz regularization using spectral normalization (SN). SN constrains the Lipschitz constant of the network by bounding the spectral norm (i.e., largest singular value) of each weight matrix, which helps stabilize training and improve robustness. As shown in the second last row in Table 5 for 4xMRI reconstruction. SN promotes adv. robustness, but at the cost of reduced accuracy to some extend. This is likely due to its restrictive nature of SN limiting the model’s expressive power.
Additionally, we also compares the performance to a diffusion models (DMs), as they also demonstrates impressive results in image generation. Since our work proposes a general framework for IPs, we chose to compare against the standard DDPM rather than specialized task-specific DMs, consistent with prior work (Güngör et al., 2023). To ensure a fair comparison under similar computational constraints, we adapt the denoising U-Net to fit within the same GPU memory as other methods.
Table 5 compares DDPM with other methods. DDPM achieves comparable in-distribution (ID) performance to MBAs trained with both MSE loss and the proposed SGD jittering, and shows stronger robustness than standard MSE-trained MBAs. However, it underperforms in OOD generalization. While DDPM serves as a strong baseline for robustness and ID accuracy, SGD jittering achieves better generalization under distribution shifts. We will add the comparison to DM in the main manuscript as an interesting baseline.
It is also worth noting that DDPM requires  10× more parameters and is significantly more data-intensive, whereas MBAs are more data-efficient (Monga et al., 2021) due to their optimization-inspired iterative structure. We also refer to prior work (Güngör et al., 2023), which compares DDPM, DiffRecon, AdaDiff to MSE-trained MBAs for MRI reconstruction. Their results show that while DM can generalize well in some cases, MBA methods consistently perform better on ID data. Results in (Güngör et al., 2023) shows that MSE-trained MBA is a strong baseline, and our proposed SGD jittering further improves their robustness and generalization.
Model Training Scheme/Design fastMRI Adv. Attack Tumor Cell
GD LU MSE training 28.21 / 0.603 25.68 / 0.382 29.92 / 0.779
AT 27.68 / 0.564 27.17 / 0.549 27.74 / 0.597
Input Jittering 28.18 / 0.595 25.05 / 0.420 29.97 / 0.740
SGD jittering (Ours) 28.22 / 0.607 26.77 / 0.552 30.36 / 0.788
Lipschitz regularization (SN) 27.71 / 0.576 27.09 / 0.542 27.92 / 0.594
DDPM Default training 28.17 / 0.611 27.13 / 0.536 29.72 / 0.782
Table 5: MRI reconstruction evaluation compared to the Lipschitz regularization method (last row). The best and second best performances are in bold and underlined respectively.
Appendix G Training Details